The following represents partial results for a one-way ANOVA: Source d.f. Sums of Squares Mean Squares F-test Between 2 300 Within Total 23 800 Given the above partially completed ANOVA table, what is the value of the F-test to determine if there is a significant difference in population means?

A. 15.67 B. 6.3 C. .375 D. 2.67
Answer #1

Solution:

We are given

MS = SS/df

F = MS between/MS within

The calculation table is given as below:

Source

df

SS

MS

F

Between

2

300

300/2 = 150

150/23.80952381 = 6.3

Within

23 – 2 = 21

800 – 300 = 500

500/21 = 23.80952381

Total

23

800

Answer: B. 6.3